OpenStudy (anonymous):

Find the diameter of the sphere. The volume is 256/3 pi yd^3

OpenStudy (jdoe0001):

$$\bf \text{volume of a sphere is}\\ V = \cfrac{4\pi r^3}{3}\\ \text{and you have }\\ V = \cfrac{256 \pi}{3}\\ \text{thus then}\\ \cfrac{256 \pi}{3}= \cfrac{4\pi r^3}{3}\\ \text{solving for "r" then}\\ \cfrac{256 \pi}{3(4 \pi)}=\cfrac{r^3}{3} \implies \cfrac{3(256 \pi)}{3(4 \pi)} = r^3\\ \cfrac{\cancel{3}(256 \cancel{\pi})}{\cancel{3}(4 \cancel{\pi})} = r^3 $$ so solving for "r", what would that give you?